### v1

# Magnetic field amplification and decay in cosmic string wakes

Abstract
We do a detailed study on vortex formation in a magnetized plasma within the spacetime of a moving cosmic string using analytical and numerical methods. The conical spacetime around the cosmic string causes the frozen-in magnetic field to deform due to the fluid flow. We find that the overdensity in the wake region amplifies the magnetic field. This amplification depends on the direction and the lengthscale of the magnetic perturbations. Alfven’s theorem of flux conservation explains this result. However, our study also shows that the magnetic field can decay depending on the perturbation lengthscale, due to the breakdown of Alfven’s theorem at a certain lengthscale. This lengthscale is the gyroradius of the charged particles in the plasma. Our findings are significant for understanding magnetic reconnection in cosmic string wakes.
journal: Eur. Phys. J. C
1 Introduction
Cosmic strings are topological defects that are formed from the symmetry-breaking phase transitions in the early universe. An interesting part of these defects is the nature of the spacetime around the string. It is known that the cosmic string metric is globally conical and has a deficit angle. Due to the nature of the metric, the geodesics of particles close to the cosmic string move inwards towards the cosmic string. This gives rise to a wake structure behind a moving cosmic string. This wake structure is well studied for an un-magnetized plasma vishniac ; sornborger ; stebbins . These wakes leave several signatures during the evolution of the universe. They can give rise to discontinuties in the Cosmic Microwave Background Radiation (CMBR)danos ; they may give rise to primordial magnetic fields vachaspati1 ; vachaspati2 ; sovan and they may also lead to density inhomogeneities which will affect the nucleosynthesis calculations layek1 ; layek ; sovan1 . Thus, understanding the structure of these wakes is important for identifying the signatures of cosmic strings in the current universe. The wake structure that has been studied in the literature is essentially the hydrodynamic wake structure. The presence of magnetic fields in the wake was generally attributed only to the superconducting cosmic strings rogozin . However, recently it has been shown that primordial magnetic fields can be also be generated in the wakes of Abelian Higgs strings sovan . So, it is possible that wakes of non-superconducting strings will also have a magnetic field. Particles moving through these magnetic fields will radiate synchrotron radiation. Such radiation can be detected by current detectors. We need to study the electron distribution, the magnetic field perturbation and many other parameters before we are able to obtain a distinct signature that points to a cosmic string. Our motivation in this study is to study in greater detail the flow of the magnetized fluid around a cosmic string and understand the important lengthscales associated with this flow.
Recent studies has shown that multiple shocks can be generated in the cosmic string wake soumen . These shocks can collide and release a large amount of energy as radiation. Moreover, recently it has also been conjectured that the narrow wake structure of the cosmic strings can lead to magnetic reconnection dilip . Lengthscales play an important role in magnetic reconnection. One of the requirements of magnetic reconnection is the breakdown of Alfven’s theorem in the magnetized plasma. This can happen at small lengthscales. We aim to identify the lengthscale at which magnetic reconnection can occur in cosmic string wakes. This lenghscale is important as it determines the energy released due to magnetic reconnection.
The fluid flow around the cosmic string is driven by the conical metric of the cosmic string. The magnetic field lines in the early universe are usually frozen into the plasma. Studies have shown that when the plasma undergoes a spherical compression, the magnetic field frozen in the plasma is also affected king . Since the flow around the cosmic string results in the fluid density increasing significantly in the wake region, we do a detailed study of the evolution of frozen-in magnetic fields around a cosmic string.
In the literature, the magnetic field in the wake region is generated due to the presence of density inhomogeneities in the plasma. These density inhomogeneities come from the clustering of different particles around the cosmic string. The clustering of these particles mostly depend on their masses abhisek . A long and narrow wake is formed. In our numerical simulations, we have found that when the velocity of the cosmic string is low, the wake structure is diffuse. A well developed wake structure is obtained for higher Mach numbers. For simplicity, we consider an ideal fluid approximation without any dissipative forces. The generated vorticity in the plasma is thus conserved. We proceed to study the evolution of the field over time in the absence of any turbulent forces.
We find that the field in the wake region can increase even without the presence of any external driving force or the presence of a dynamo mechanism. This is expected as the field lines are frozen in and so the conservation of magnetic flux density usually enhances the field when the fluid flow squeezes the flux lines. So, it is not surprising that the magnetic field is amplified due to the overdensity of the wake. However, we also find that in certain cases, the magnetic field decreases rapidly in certain directions, which demonstrates that not all magnetic fields get amplified in the cosmic string wake. We find that the amplification or decay depends on the length scale of the magnetic field perturbation. A detailed analysis reveals that the decay of the magnetic field is due to the breakdown of Alfven’s theorem at very small lengthscales. In the numerical set up, we use a exponentially decreasing magnetic field to illustrate our findings.
Our detailed study also shows that if the sheared magnetic field perturbation in the cosmic string wake is aligned to the flow of the plasma past the string, then the magnetic field is always amplified for whatever perturbation is given. It is only when the magnetic field perturbation is perpendicular to the direction of the flow, the decay lengthscale of the perturbed magnetic field determines whether the field gets amplified or whether it decays. This is because the frozen in magnetic field is affected more by the perpendicular component of the velocity of the fluid rather than the parallel component. We have presented results for a cosmic string moving in a high $\beta$ plasma as the wake formation in our numerical study is more pronounced for the high $\beta$ plasmas. For the numerical part, we have used the OpenMHD code, an open source code, which is available from GitHub.
In section II, we discuss the spacetime around a cosmic string and discuss how vorticity is generated in a cosmic string wake. In section III, we discuss the amplification of the magnetic field due to the deformation of the magnetic flux lines. In Section IV, we show the decay of the magnetic field in presence of a sheared magnetic field. We also show that in the presence of a sheared magnetic field, the amplification or decay depends on the orientation and the lengthscale of the perturbative field. Section V we summarize the work and discuss why it is important to understand the evolution of magnetic fields in the cosmic string wakes for current observational signatures.
2 Vorticity in a cosmic string wake
The geodesic equations around a cosmic string metric have been studied in detail in the literature abhisek . It has been shown that particles can have closed or open orbits around an Abelian Higgs string based on their total angular momentum and energy. In general, particles moving past a cosmic string suffer a change in their velocities close to the cosmic string. This leads to the wake formation behind moving cosmic strings. The metric being symmetric around the deficit angle of the string, the wake structure is also symmetric around the deficit angle of the string.
The metric of a cosmic string is given by vilenkin ; deruelle ,
$ds^{2}=N^{2}(\rho)dt^{2}-d\rho^{2}-L^{2}(\rho)d\theta^{2}-N^{2}(\rho)dz^{2}$ (1)
The factors $L(\rho)$ and $N(\rho)$ are determined by the boundary conditions. They are related to the values of the fields at a distance $\rho$. The angle $\theta$ varies from 00 to $(1-4G\mu)\times 2\pi$. The missing angle $\Delta\theta=8\pi G\mu$ is referred to as the deficit angle of the cosmic string.
Apart from the presence of the deficit angle, the metric of the cosmic string is locally flat. So, the ideal magnetohydrodynamic (MHD) equations remain the same. They are given by mkverma ,
$\frac{\partial\rho}{\partial t}+\nabla.(\rho\mathbf{v})=0$ (2)
$\rho\left(\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}.\nabla)\mathbf{v}\right)=-\nabla\left(P_{th}+\frac{B^{2}}{8\pi}\right)+(\mathbf{B}.\nabla)\mathbf{B}$ (3)
$\frac{\partial\mathbf{B}}{\partial t}+(\mathbf{v}.\nabla)\mathbf{B}=(\mathbf{B}.\nabla)\mathbf{v}$ (4)
The only difference comes as the particles move past the string. The conical nature of the metric gives a velocity perturbation to the particle towards the cosmic string. This velocity perturbation is proportional to the deficit angle $\Delta\theta$ of the string. It is given by $\delta v\approx\delta\theta v_{s}\gamma_{s}$, where, $v_{s}$ is the cosmic string velocity and $\gamma_{s}$ is the Lorentz gamma factor corresponding to the moving cosmic string. This causes the particle trajectories to bend towards the moving cosmic string. Though small, $\delta\theta$ is always non-zero, hence, ×$\triangledown\times\mathbf{v}$ is non-zero. A rotational flow is thus generated. The local vorticity of the plasma is defined by,
$\boldsymbol{\omega_{L}}=\triangledown\times\mathbf{v}$ (5)
Here, $\mathbf{v}$ is the particle velocity. Since the deficit angle is small, the vorticity is also generated at smaller lengthscales.
In the case of magnetohydrodynamics, the vorticity may be enhanced by the presence of the charged paricles in the plasma. For the ideal fluid approximation, the magnetic field lines are frozen into the flowing fluid. It is possible to define a generalized vorticity for the fluid which considers the magnetic field. Charged particles move in rotational motion in a magnetic field, hence the vorticity is enhanced. The generalized vorticity is defined by mahajan ,
$\boldsymbol{\omega_{G}}=\triangledown\times\mathbf{v}+\frac{q\mathbf{B}}{c}$ (6)
For an ideal fluid, the vorticity generated in the fluid will always be conserved. In most systems with ideal fluids, it is the boundary conditions which lead to the generation of vorticity but in the case of the particles moving close to a cosmic string, an angular momentum is generated as the particles orbit the string. The orbital angular momentum then leads to a non-zero ×$\triangledown\times\mathbf{v}$ in the fluid. So, the vorticity is generated in the wake of a cosmic string as it moves through the cosmic fluid. In the case of hydrodynamics, this local vorticity is small but in the case of magnetohydrodynamics, this vorticity is important for identifying lengthscales of the magnetic field perturbations. We show later on why “coarse-graining” the fields and the smallest lengthscale are important for the growth or decay of the magnetic field.
2.1 Numerical Simulation of the vorticity
We use the publicly available OpenMHD code openmhd to simulate the vorticity in the cosmic string wakes. We use the basic code and use open boundary conditions in the direction of fluid flow. In the basic code, the MHD equations along with the equation of state, relating the hydrodynamic pressure and the density, $P_{H}=f(\rho)$ are solved using a perturbative method. For the case of magnetogydrodynamics, the total pressure ($P$) consists of the hydrodynamic pressure and the pressure due to the magnetic field mkverma . The fields ($\mathbf{v}$, $\mathbf{B}$, $P$, $\rho$) mentioned previously in the MHD equations can be split into a background field and a perturbative field as, $\mathbf{v}=\mathbf{v_{0}}+\mathbf{v_{1}}$, $\mathbf{B}=\mathbf{B_{0}}+\mathbf{B_{1}}$, $P=P_{0}+P_{1}$ and $\rho=\rho_{0}+\rho_{1}$. Here $(\mathbf{v_{0}}$, $\mathbf{B_{0}}$, $P_{0}$, $\rho_{0})$ are the background fields, and $(\mathbf{v_{1}}$, $\mathbf{B_{1}}$, $P_{1}$, $\rho_{1})$ are the perturbative fields. A Galilean transformation is used to shift to the background flow frame, thus, $\mathbf{v_{0}}=0$, $\mathbf{v_{1}}=\mathbf{v}$ which means that the velocity field $\mathbf{v}$ is the perturbative field. These are then substituted in the MHD equations mentioned previously, and only the first order terms are retained, assuming the perturbations to be small. The linearized equations are given by,
$\frac{\partial\rho_{1}}{\partial t}+\nabla.(\rho_{0}\mathbf{v})=0$ (7)
$\rho_{0}\left(\frac{\partial\mathbf{v}}{\partial t}-(\mathbf{B_{0}}.\nabla)\mathbf{B_{1}}\right)=-\nabla P_{1}-\rho_{0}\nabla(\mathbf{B_{0}}.\mathbf{B_{1}})$ (8)
$\frac{\partial\mathbf{B_{1}}}{\partial t}-(\mathbf{B_{0}}.\nabla)\mathbf{v}=-\mathbf{B_{0}}\nabla.\mathbf{v}$ (9)
The basic code is available in one and two dimensions. We have used the 2D code and modified it by giving a velocity perturbation to the particles as they cross the cosmic string. The fluid flow is taken to be along the $x$-axis. In the $y$-direction, we have used fixed boundary conditions. The cosmic string is aligned with the $z$-axis, it is identified in the flow by the position at which an initial velocity perturbation is given to the plasma particles. All the particles are subject to this velocity perturbation which will depend on the deficit angle of the cosmic string metric. Thus, at the initial time, plasma particles crossing a certain position ($x$ coordinate) are all given a velocity perturbation both in the $x$-and $y$-direction; $v_{x}=-v_{0}Cos\theta$ and $v_{y}=-v_{0}Sin\theta$. All other particles move with velocity $-v_{0}$. This is in the reference frame of the cosmic string. As the particles flow, after some time the vortex structure is formed. Fig.1 shows a streamline plot of the vortex structure generated in a cosmic string wake.
Figure 1: Generation of a pair of vortex in the cosmic string wake.
For a better understanding of the simulation, a larger deficit angle is taken so that the vorticity is clearly visible. However, it is possible to scale the results to the cosmic string lengthscales. As mentioned before, the deficit angle for an actual cosmic string is very small, in dimensionless parameters the perturbation would be of the order of the deficit angle with a value of $10^{-5}$ gangui . Based on this, the lengthscale of the vortex will be of the order of $10^{-5}Mpc$ or smaller. The net vorticity over a certain region is calculated and in the absence of a magnetic field, we find that the vorticity is conserved throughout the flow. This is an important check, as we have mentioned that we do not consider any dissipative forces in the system, so the energy conservation will also mean the conservation of vorticity in the flow.
2.2 Magnetic fields in rotational flows
We now consider the role of the magnetic field in the simulation. The magnetic field $B$ has two components (in $x$-and $y$-direction). For a uniform magnetic field, we give fixed values to both components. Generally, the generation of vorticity in magnetohydrodynamics leads to turbulence in the plasma. However, for a turbulence to develop, we need to introduce a constant source for generating vortices at different length scales. As mentioned before, we have a constant background magnetic field in the simulation so the vorticity is generated at a fixed lengthscale. For turbulence, the plasma has to be non-ideal and the net vorticity need not be conserved. In our simulation, we do not introduce any other potential so the lengthscales of the generated vortices are fixed by the scale of the cosmic string wake and the background magnetic field. Hence no turbulence occurs in our system. So, the dynamo mechanism is completely absent from our simulations.
We plot the magnetic field lines in the simulation. As expected, the magnetic field lines are deformed due to the flow of the fluid. We calculate the total magnetic field in the wake region as well as the peak value of the field. As long as the magnetic field remains constant, we do not expect any change in the magnetohydrodynamic flow. The flow remains steady with minor changes due to the velocity perturbation. We proceed to give a perturbation to the magnetic field. Since we would like the field to be amplified, we give a sheared perturbation to the magnetic field as it is previously known that sheared magnetic fields may lead to amplification of the magnetic field king . The field is perturbed first along the $x$- direction and then in the $y$-direction, the perturbations are denoted by $B_{x}$ and $B_{y}$, respectively. The results of these perturbations are discussed in the next section.
3 Sheared magnetic fields and field amplification
For a frozen in magnetic field, the frozen in condition ensures that the magnitude of the magnetic field will be related to the density of the plasma. It has been shown that an isotropic collapsing plasma can lead to the amplification of a magnetic field even in the absence of a dynamo mechanism king . The argument was that the flux in the frozen in plasma is conserved, so in the collapsing region, as the cross-sectional area decreases, the magnitude of the magnetic field increases. Now, the formation of a wake behind a cosmic string leads to a substantial increase in the density of plasma behind the cosmic string. In this case too, there is the possibility that the magnitude of the magnetic field will be enhanced.
To understand this, we looked at two different kinds of magnetic field perturbations. We start with an exponentially decreasing magnetic field perturbation. The strength of the magnetic field is proportional to the density of the magnetic field lines. As mentioned before, the cosmic string is moving in the $x$-direction. Due to the velocity perturbation, the fluid velocity has a component in the $y$-direction, and therefore, the magnetic field lines feel a force along the $y$-direction. Therefore, we start with a sheared perturbation in the direction of motion of the cosmic string. Fig.2 represents the graph of the total magnetic field with time for a perturbation of $B_{x}=B_{0}e^{-\alpha_{1}|y|}$. We call this a sheared magnetic field as the $x$ component of the magnetic field ($B_{x}$) at any given point has a variation in the $y$-direction. The lengthscale of the variation is given by the parameter $\alpha_{1}$. We have used $B_{0}=1$.
Figure 2: Evolution of the magnetic field energy in the cosmic string wake for a value $\alpha_{1}=0.01$. The magnetic field energy is given in scaled units and hence is dimensionless
Figure 3: Evolution of the peak value of the magnetic field in the cosmic string wake. The value after a few time steps is five times the initial value of the magnetic field.The magnetic field is given in scaled units and hence is dimensionless.
Fig.3 represents the graph of the peak magnetic field evolution with time for $\alpha_{1}=0.01$. As the graphs shows there is an increase in the peak magnetic field value which will lead to the total magnetic energy being enhanced. The figure shows that the peak magnetic field value becomes at least five times the initial value. The peak magnetic value shown on the $y$- axis is the maximum value of the magnetic field at a specific point on the grid. Similarly, the total energy also becomes five times the initial energy value, the $|B|^{2}$ on the y - axis is the summation of the magnetic energy density over all the grid points. Together, they indicate that the magnetic field value not only increases at a particular point, it increases over the whole of the cosmic string wake. So, we see that the net magnetic field gets amplified without any external dynamo mechanism. We now discuss in detail why this is happening.
The velocity perturbation on the particles is along the $y$-direction, so the particles move towards the conical space behind the cosmic string. We have given a sheared magnetic field as the magnetic field perturbation. Though we are perturbing the $x$ component, the variation is along the $y$-direction. As the magnetic field lines are being pushed in the $y$-direction by the velocity flow, so our perturbation is enhancing the effect of the compression of the magnetic field lines due to the fluid flow. We find that we get an amplification of the field for all values of $\alpha_{1}$.
The amplification of the magnetic field without any external turbulent forces can be attributed to the deformation tensor of the fluid element. The deformation tensor of the fluid element has been used to explain the amplification of frozen in magnetic fields in ref king . The deformation parameter $D_{ij}$ represents the deformation of an infinitesimal fluid element of initial side length $\delta q_{j}$, which in this case is deformed due to the deficit angle in the local space of the cosmic string. The deformation tensor is defined by
$D_{ij}\equiv\frac{\partial x_{i}}{\partial q_{j}}.$ (10)
Fig.4 is used to explain how the magnetic field gets amplified for this given magnetic field perturbation. As has already been pointed out in ref king , the magnitude of the magnetic field is affected only when the fluid motion is perpendicular to the magnetic field lines. In this case, for any given value of $x$, the magnetic field varies along $y$ as an exponentially decaying field. This is illustrated in Fig.4. The $q_{1}$ in the figure is the $x$-axis, while the $q_{2}$ is the $y$-axis. The angle $\theta$ by which the fluid element is deformed is proportional to the deficit angle of the cosmic string. The comoving volume element is preserved after the deformation, so $\frac{V^{fin}}{V^{ini}}=D_{ij}$. Since the volume can be related to the density of the fluid, we will have
$\frac{V^{fin}}{V^{ini}}=\frac{\rho^{ini}}{\rho^{fin}}=D_{ij}$ (11)
As we are considering frozen in magnetic fields, squeezing or stretching of the field lines will depend on the change in the cross-sectional area. Therefore, we will also have,
$\frac{B^{fin}}{B^{ini}}=\frac{A^{ini}}{A^{fin}}$ (12)
where $A$ denotes the initial and final cross-sectional areas through which the magnetic field line passes before and after deformation. If the length perpendicular to this area is referred to as $L$, from volume conservation we will be have,
$\frac{B^{fin}}{B^{ini}}=\frac{L^{fin}}{L^{ini}}$ (13)
The angle between $dq_{2}$ and $L^{fin}$ is the deficit angle $\theta$. This vector $L^{fin}$ has not been shown on the Fig.4 as it is rather difficult to depict in this case. Thus we will have
$\frac{B^{fin}}{B^{ini}}=\frac{\rho^{fin}}{\rho^{ini}}\frac{L^{ini}}{L^{fin}}$ (14)
In the case, where $\frac{L^{ini}}{L^{fin}}\approx 1$, the final magnetic field will be determined by the overdensity of the wake.
Figure 4: The figure representing the deformation of the magnetic field lines in a fluid flowing around a cosmic string. The point at which the axes $dq_{1}$ and $dq_{2}$ meet is the place at which the cosmic string is placed.
In this case, the perturbative field is varying in the $y$-direction. So, we also have off-diagonal elements which are non-zero. The off-diagonal elements correspond to terms like $\frac{\partial x_{2}}{\partial q_{1}}$. Here, these terms are non-zero as $\partial x_{2}$ refers to the change in the $y$-direction corresponding to a small change in the $x$-direction. In all the cases that have been discussed in ref. king , the anisotropic collapse has lead to the magnetic field being proportional to the density of the fluid. As the wake has reflection symmetry about the $x$-axis, we will have $D_{12}=D_{21}$, and the other off-diagonal elements can be treated as zero. In such a case, we have
$\frac{B^{fin}}{B^{ini}}=\frac{1}{(1+D_{12})^{1/2}}$ (15)
This will mean that $B\propto\rho^{1/2}$. Similarly in any case, that is considered, the final magnetic field is determined by the final density of the deformed fluid.
Now, in the wake region, we will always have $\rho^{fin}>\rho^{ini}$, therefore the magnetic field will always be amplified in the wake of the cosmic string due to the deformation of the field lines in the presence of the deficit angle. This will happen as long as the magnetic field remains frozen in the plasma. We have checked the results for different values of $\alpha_{1}$ and in all the cases we see an increase in the magnitude of the magnetic field in the wake region.
The magnetic field in the cosmic string wake can have components in both the $x$-and $y$-direction. So, our next case would be a perturbation of $B_{y}=B_{0}e^{-\alpha_{2}|(x-800)|}$. The origin is shifted to get a better figure of the final wake. It does not have anything to do with the field magnitude. Here, we depict the lengthscale by the parameter $\alpha_{2}$. The initial value is still $B_{0}=1$.
Figure 5: Evolution of the magnetic field energy in the cosmic string wake for a value $\alpha_{2}=0.01$.
Figure 6: Evolution of the peak value of the magnetic field in the cosmic string wake. The value satuarates to about five times the initial value even in this case.
As is seen from Fig.5 and Fig.6, we still get a amplification of the magnetic field. The maximum value is approximately five times the initial value again. In this case too, the field lines are deformed by the perturbation. As we have taken the flow velocities and the perturbations along the $x$-and $y$-axes, the amplification is similar in both the cases. If the axes were not aligned we might have got different results for the amplification. In both our cases the off-diagonal elements which are non-zero are driven by symmetry consideration. So, we can show that the magnetic field is proportional to the density of the wake in both the cases. We have repeated the simulation with a different periodic magnetic field and obtained similar results about the magnetic field amplification. For simplicity, we have taken one of the axes to be aligned with the flow, even if that is not the case, the amplification will still happen, only the symmetry consideration cannot be used to calculate the off-diagonal elements.
4 Magnetic field decay
While surveying the various values of $\alpha_{1}$ and $\alpha_{2}$, we also found that while the field always increases for any value of $\alpha_{1}$, this is not the case for $\alpha_{2}$. For $\alpha_{2}=1$, we find that the field decays with time. This is plotted in Fig.7 and Fig.8.
Figure 7: Evolution of the magnetic field energy in the cosmic string wake for a value $\alpha_{2}=1$.
Figure 8: Evolution of the peak value of the magnetic field in the cosmic string wake at a particular point in the wake.
So, for a perturbation perpendicular to the flow velocity of the plasma, the lengthscale of the perturbation determines whether the field decays or it gets amplified. As discussed in the previous section, the Alfven’s theorem leads to the frozen in magnetic field lines. Due to the deformation of the fluid, the magnetic field lines get squeezed, and the amplification of the field is observed. However, in this case too the fluid is deformed but the magnetic field does not get amplified. This seems to indicate that Alfven’s theorem is no longer valid for this lengthscale.
The breakdown of Alfven’s theorem has been studied in connection to ideal plasma flows eyink . This breakdown leads to the phenomenon of magnetic reconnection in space plasmas. It has been shown that flux conservation is violated at an arbitrarily small lengthscale ($l_{0}$) for certain conditions. For such lengthscales, dissipative non-ideal effects will be significant. The question arises how do we identify this lengthscale in our problem. Our numerical simulations indicate that it is related to the value of $\alpha_{2}$ of the perturbative magnetic field. So, we concentrate on the lengthscale $l\sim 1/\alpha_{2}$, which we denote by $l_{0}$. The magnetic field can be considered to be “coarse-grained” at lengthscale $l_{0}$ so that at scales smaller than this lengthscale the fluid becomes non-ideal.
In the fluid flow of the wake, the magnetic field lines initially move with the velocity of the fluid. The advection velocity of the loops are then given by the velocity of the particles in the plasma. Assuming the majority of the particles to be electrons, one can then calculate the gyroradius of the plasma electrons. The gyroradius is given by $r_{g}=\frac{mcv_{y}}{qB}$ where $v_{y}$ is the perpendicular component of the velocity of the electron. We find that the gyroradius gives us a good estimate of whether $l_{0}$ has been reached. Based on the fluid velocity, we calculate an approximate value of the electron gyroradius $r_{g}$. We find that the magnetic field gets amplified only when the particle gyroradius $r_{g}$ is smaller than the scale at which the magnetic field perturbation varies, i.e. for $r_{g}<1/\alpha_{2}$. Since in our numerical simulations the value of $r_{g}\sim 1.31$, hence we find that when the value of $\alpha_{2}\sim 1$, the field decays instead of getting amplified.
The relation of the lengthscale $l_{0}$ to the gyroradius can be obtained from the trajectory of charged particles in non-uniform magnetic fields. Previously, the motion of charged particles has been studied in specific sheared force-free magnetic fields vekstein . In this case, we have a non-uniform sheared magnetic field perturbation perpendicular to the flow direction of the fluid. The motion of the electrically charged particles in a perturbed magnetic field has been studied using the approximate description called the guiding center or drift approximation northrop . The charged particle motion in the plasma is decomposed into a nearly circular fast gyromotion about a guiding center and the trajectory of the guiding center itself. Neither of the trajectories are constant for the non-uniform magnetic field. The overall motion is obtained by solving the second-order differential equations for the guiding center. Usually, for better understanding, the equations are separated into parallel and perpendicular components of the given magnetic field. In our present case, the perpendicular direction is the direction along which the field is non-uniform. The velocity in the perpendicular direction (i.e. along the non-uniform direction) is then given by a gradient drift ($v_{g}$) and a curvature drift ($v_{c}$). These are defined by,
$\boldsymbol{v_{g}}=\frac{mv_{y}^{2}c}{2qB}\left(\mathbf{h}\times\frac{\nabla\mathbf{B}}{B}\right)$ (16)
$\boldsymbol{v_{c}}=\frac{mv_{x}^{2}c}{qBR^{2}}\left(\mathbf{R}\times\mathbf{h}\right)$ (17)
Here, $\mathbf{h}=\frac{\mathbf{B}}{B}$ is the unit vector along the magnetic field line and $\mathbf{R}$ is its radius of curvature. From the definition of the gyroradius and the fact that the term $\frac{\nabla\mathbf{B}}{B}$ will be inversely proportional to the lengthscale of the varying magnetic field, we can estimate the order of magnitude of the drift velocity of the particles will be given by $v\times\frac{r_{g}}{L}$. Here, $L$ is the spatial scale of the non-uniformity of the magnetic field. For $r_{g}<<L$, the drift velocity approximation is valid. Since this approximation indicates that the velocity of the magnetic lines can be looked as the velocity of the trajectory of the center of the gyro motion, Alfven’s theorem will be valid in this approximation. For $r_{g}\sim L$, the lengthscale of the variation of the magnetic field will be the same as the gyroradius, so the gyroradius cannot be defined at all for these electrons. This is because the underlying assumption in defining the gyroradius is that the field should be uniform. So, the concept of the guiding center cannot be used to understand the movement of the magnetic field lines. So, for $r_{g}\geq L$, the magnetic field lines no longer behave as if they are frozen in the plasma.
Thus, we conclude that sheared magnetic fields with lengthscales greater than the gyroradius will be amplified in the cosmic string wake, whereas magnetic fields with smaller lengthscales will decay. So, for sheared perturbations of small lengthscales, Alfven’s theorem is no longer valid. This is similar to the situation in the Sweet-Parker model of magnetic reconnection. In that case, two opposite fields close to each other generate a neutral region where the magnetic lines of force can break and reconnect. This is the resistive diffusive region; outside this region, Alfven’s theorem of frozen in magnetic field lines still holds while inside this region, the magnetic lines of force break and form an X-type neutral point reconnection . Since it has been demonstrated that magnetic reconnection can happen in cosmic string wakes dilip , it is important to understand the small-scale dynamics of the magnetic field in the cosmic string wakes. From this study, it appears that the lengthscale of the reconnection region for cosmic string wakes should be less than the gyro radius corresponding to the magnetic field in the wake.
5 Summary and Conclusions
We have done a detailed study of the amplification of magnetic fields in the wake of a cosmic string in the absence of a dynamo mechanism. Though hydrodynamic wakes and shocks are well studied for a cosmic string, it is only in recent times that magnetohydrodynamic shocks have also been studied in the cosmic string wake. Magnetohydrodynamic shocks are difficult to study as the magnetic field lengthscales play a major role in the generation and evolution of the shock dahlburg . Among cosmic string wakes, MHD shocks have been discussed previously in ref. beresnyak . In recent times, more detailed simulations of such wakes have been carried out in ref. soumen . One of the crucial factors that were never considered in these studies was the lengthscale of the perturbations that affect the magnetic and density fields. Lengthscales of both density inhomogeneities as well as magnetic fields affect structure formation and later evolution of large scale structures in the universe kim . For cosmic strings, this is particularly important as density perturbations depend on the small-scale structure of the strings wu as well as the nature of the particle motion around them abhisek .
In this respect, our study is important as it clearly shows the effect of perturbations of different lengthscales on the magnetic field. We find that the magnetic field gets amplified by the presence of sheared perturbations close to the core of the cosmic string. We also find that the amplification strongly depends on the nature of the magnetic field and it’s lengthscale. The amplification of the magnetic field occurs due to the squeezing of the magnetic flux in the wake of the cosmic string. This can only happen when the magnetic field lines are frozen into the plasma. The shear causes a deformation in the fluid. Assuming the wake is symmetric about the flow axis, we show that the deformation tensor will have non-zero off-diagonal terms. This leads to the fact that the magnetic field strength will become proportional to the square root of the fluid density. Depending on the orientation of the deformation, the magnetic field can also be proportional to the density. Since the overdensity in the wake is always greater than the initial density of the flow, the magnetic field flux gets amplified in all the different cases.
We then proceed to show that the amplification also depends on the nature of the perturbation of the magnetic field. For a sheared perturbation which is perpendicular to the direction of flow of the fluid, the magnetic field lengthscale determines whether the magnetic field will amplify or decay. This is because below a certain lengthscale the Alfven’s theorem of frozen in magnetic field may not be valid and local interactions will determine the dissipation of the magnetic field. We have determined that the length scale below which the magnetic field decays is determined by the gyroradius of the particles being considered. If the lengthscale of the magnetic field is much larger than the gyroradius, Alfven’s theorem remains valid and we get the amplification of the field. However, if the lengthscale of the magnetic field perturbation is of the same order or less than the gyroradius of the electron then the field decays.
The magnetic field due to the Biermann mechanism is due to the small-scale inhomogeneities generated in the cosmic string wake. Since the density perturbations of a long cosmic string are of varied lengthscales based on how they are generated, it is important to understand the behaviour of the magnetic field for perturbations of different lengthscales. There are cosmic strings with small-scale structures on them which can generate small-scale non-linear density perturbations. It is also known that non-Gaussian density perturbations can be generated by long strings dacunha . As long as Alfven’s theorem holds, the magnetic field configurations will directly relate to the density perturbations. In this work, we have shown that the lengthscale of the magnetic field perturbations will determine if Alfven’s theorem is applicable and we will have a magnification of the magnetic field. We have also seen that for magnetic perturbations with lengthscales of the order of the gyro radius or less, Alfven’s theorem breaks down. This in turn can lead to the possibility of magnetic reconnection in the cosmic string wake. Since, in recent times, magnetic reconnection has been seen as a viable possibility in the narrow wakes of the cosmic string, this study will lead to a further understanding of the evolution of magnetic fields in cosmic string wakes.
Our study has some limitations as it is constrained to two dimensions only. The current study therefore has one dimensional sheared magnetic fields only. The actual wake structure is three-dimensional, so it is also possible that the perturbation is a planar sheared field. The evolution of such perturbations may lead to further interesting results. We have also considered ideal MHD only, it is quite possible that the reconnection possibility and the breakdown of Alfven’s theorem will be different for resistive MHD. In fact, there might be a higher probability of obtaining magnetic reconnection in that case. These and other aspects of magnetic field evolution in cosmic string wakes need to be studied in more details in the future.
Acknowledgements.
We acknowledge funding from the DST-SERB Power Grant no. SPG/2021/002228 of the Government of India. D.B was supported through the Science Academies Summer Research Fellowship program 2023, D.K is supported by the Power Grant no SPG/ 2021.002228, and S.N acknowledges financial support from the CSIR fellowship no 09/414(2001)/2019-EMR-1 given by the HRD, Govt. of India.